# Reconstructing the unitary part of a noisy quantum channel

### v2

@@ Abstract
We consider the problem of reconstructing, from a set of input and output states, the unitary describing the evolution of a quantum system or quantum channel. For ideal, fully coherent evolution, we show that the unitary can be reconstructed from two mixed states or $d+1$ pure states, where $d$ is the size of Hilbert space. The reconstruction method can be extended to approximate the unitary part of a dynamical map, provided the decoherence is not too strong to render this question meaningless. We exemplify the method for the example of the cross-resonance gate as well as a random set of unitaries, comparing the reconstruction from pure, respectively mixed, states to an approach based on the Choi matrix. We find that the pure state reconstruction requires the least amount of resources when the dynamics is close to unitary, whereas the mixed state approach outperforms the pure state reconstruction in terms of channel uses for appreciable decoherence.
@@ I Introduction
Here, we propose an alternative approach. The setup we envision is as follows: Considering an unknown or badly characterized quantum device, which input states should be prepared, sent through the channel, and measured on output, in order to minimize the number of channel uses? For an ideal, i.e., unitary, channel, the task of identifying the unitary can be thought of as two subtasks — first, to identify the basis in which the unitary is diagonal and then to determine the eigenvalues via their phases. The minimum number of states is therefore at least two, one that is built up of one-dimensional projectors, weighted with mutually disjoint eigenvalues, allowing to assess the information on the basis, and a second one for the phases, that is rotated with respect to all the one-dimensional projectors [5]. The second state can be thought of as a basis state taken from a mutually unbiased basis. The two states can be used to estimate the average fidelity of a quantum gate [5, 6]. Here we show that measuring these two states upon output allows not just to estimate the gate fidelity, but also to reconstruct the unitary part of the channel, provided that the decoherence is not too strong. Instead of identifying the basis via a single mixed state, it is also possible to directly employ $d$ one-dimensional projectors. This ensures that all input states are pure but takes their from two (irrespective of Hilbert space dimension) to $d+1$, with $d$ the Hilbert space dimension. We compare our state-based reconstruction approaches to a reconstruction using the Choi matrix.
The paper is organized as follows: In Sec. II we present our two state-based reconstruction approaches for a unitary quantum channel using the minimal state set involving mixed, respectively pure, states. The generalization to noisy quantum channels is discussed in Sec. III, where we also outline the reconstruction based on the Choi matrix, as a reference point for our later analysis. In Sec. IV we apply the reconstruction algorithms via numerical simulations to a physical model involving typical relaxation processes. We start with two-qubit examples and then investigate the extension to multiple qubits. In Sec. V we compare the practical resource cost of the reconstruction algorithms in the context of state-of-the-art compressed sensing approaches for the two cases of the channel being close to unitarity and for appreciable channel noise. Section VI summarizes our findings.
@@ II Reconstruction Approaches for Unitary Dynamical Maps
The action of a dynamical map on only two density matrices is sufficient to distinguish any two unitaries or to assess whether a given dynamical map is unitary [5]. Alternatively, it is possible, to employ a set of $d+1$ pure states, where $d$ is the dimension of the underlying Hilbert space. Accordingly, the action of a dynamical map on at least two density matrices, respectively $d+1$ pure states, is required to reconstruct the unitary operation $U$ describing the action of a unitary dynamical map $\mathcal{D}_{U}(\rho)=U\rho U^{\dagger}$. In this section, we explain how the unitary $U$ can be obtained from the action of the dynamical map on the two minimal state sets mentioned above, making no further assumptions on the dynamical map except for its unitarity.
@@ IV Application
Next we test and compare the reconstruction methods outlined in Sec. III. We begin by introducing the physical model in Sec. IV.1 and introduce the gate error $1-F_{G}$ in Sec. IV.2 as the main figure of merit to quantify the performance of the reconstruction methods. After analyzing the case of two qubits in Sec. IV.3, we investigate how the performance of the methods scales with the number of qubits in Sec. IV.4.
@@ IV.2 Figure of merit
The variance in Eq. (19) is taken over unitary maps with corresponding Choi matrix $\chi^{k}_{CL}$ corresponding to a Clifford gate $U^{k}_{CL}$ randomly drawn from the Clifford group. We use XXX random Clifford gates to evaluate the variance of their average fidelites with respect to the Choi matrix.
@@ IV.3 Results: Two qubits
Figure 1: Gate error $1-F_{G}(U_{0},U_{RC})$ versus $T_{1}$ (left y-axis) of the reconstruction methods for random two qubit gates (left) and a CNOT gate (right). The shaded areas (left) represent the $1\sigma$ deviation with respect to the set of 100 random unitary gates drawn according to the Haar measure. The plot also shows the unitarity error $1-u(\chi)$ (right y-axis). The dissipation on the qubits is given by $L_{1}=\frac{1}{T_{1}}\sigma^{-}\otimes\mathds{1}$ and $L_{2}=\frac{1}{T_{1}}\mathds{1}\otimes\sigma^{-}$. The dissipation strength $T_{1}$ is given in units of the gate time $T_{g}$.
Figure 1 shows the gate error $1-F_{G}$ for the mixed-state, pure-state, and Choi reconstruction method versus the $T_{1}$ decay time for random two-qubit operations (top panel) as well as the CNOT gate (bottom panel). It also shows the unitarity error $1-u(\chi)$ of the Choi matrix $\chi$ corresponding to the full dynamical map $\mathcal{D}(\cdot)$ according to Eq. (11). Note that the dissipation strength $\gamma_{1}$ of the dynamical map increases from right to left on the x-axis due to $\gamma_{1}=\frac{1}{T_{1}}$. Figure 1 clearly demonstrates that the pure-state and mixed-state reconstruction methods can be employed for dynamics that are sufficiently close to unitary, cf. the unitarity error (red line). Even when the decay time $T_{1}$ decreases, i.e. the dissipation strength increases, and the unitarity error of the dynamical map becomes larger, the two methods are able to approximately reconstruct the unitary part of the dynamical map for a wide range of dissipation strengths.
The lines for the gate error of the mixed-state and pure-state reconstruction do not cover the entire range of $T_{1}$ times. This is due to the breakdown of the methods for large deviations from unitarity: When the dissipation strength increases, the assumptions used in Sec. II are no longer fulfilled, and degeneracies in the spectrum of the images $\mathcal{D(\rho)}$ prohibit the use of the pure-state and mixed-state reconstruction methods. In contrast, the reconstruction via the Choi matrix is not influenced by degeneracies in the image. Hence the corresponding gate errors stretch across the entire range of $T_{1}$ times. Figure 1 furthermore shows that, for the mixed-state reconstruction, degeneracies in the image of the dynamical map occur for smaller dissipation strengths compared to pure-state reconstruction (compare the $T_{1}$-coverage of the orange and blue lines in Fig. 1). This is because the pure-state images in Eq. (7) are more protected against degeneracies than the mixed-state image (see Eq. (3)). It reflects the fact that for pure initial states it is less likely that the dissipative part of the dynamical map annuls the spectral gap between eigenvalues of different eigenvectors than it is for mixed initial states.
@@
We observe in Fig. 1 that the mixed-state reconstruction performs worse in terms of the gate error $1-F_{G}$ than the pure-state reconstruction, while the Choi reconstruction performs best. This behavior is intuitive if one considers the total amount of information available: The Choi matrix contains, via Eq. (9), the information of the action of the dynamical map on arbitrary input states. The total amount of information accessible to the algorithm for the reconstruction is therefore higher compared to mixed-state and pure-state reconstruction. Similarly, since for mixed-state reconstruction a smaller number of states is propagated, less information is gained compared to the pure-state reconstruction. This explains the difference in gate errors between mixed-state and pure-state reconstruction.
@@
For the mixed-state reconstruction of random gates, there is no appreciable difference in the general behavior as a function of $T_{1}$ and $T_{2}$, cf. the top panels of Fig. 1 and Fig. 2. In particular, larger gate errors compared to the pure-state reconstruction method are observed for both dissipation processes. For the CNOT gates, two observations can be made: The gate errors for both mixed-and pure-state reconstruction are larger than for the random gates. The dissipation strength, at which the image becomes degenerate, is larger for pure dephasing compared to $T_{1}$ decay (cf. the ranges of the blue and orange lines in the lower panels of Fig. 1 and Fig. 2). This is in contrast to the random gates, where there is no significant difference between $T_{1}$ and $T_{2}$ for both mixed-state and pure-state approach.
@@ IV.4 Results: Multiple qubits
How does the system size affect the performance of the reconstruction methods? To answer this question, we consider random unitaries on a system of $N$ qubits for the case of $T_{1}$ decay. We have seen in Sec. IV.3 that the gate error of the reconstruction methods is connected to the unitarity of the dynamical map. The specific form of the dissipation determines unitarity, and we consider two choices in the following. In the first choice, all qubits are subject to decay, with identical dissipation rates $\gamma_{1}^{(i)}=\frac{1}{T_{1}}$ for $i=1,\dots,N$. Such a scenario is common in qubit arrays where the error is caused by spontaneous decay [16]. In this case, one would expect the gate error to increase with system size, since the total dissipation strength is $\gamma_{1}^{\text{tot}}=\sum_{i}^{N}\gamma_{1}^{(i)}$ which scales with $N$. Figure 3 (top) shows the gate error versus $T_{1}$ times for random unitaries for $N=2,3,4,5$ qubits, reconstructed using the mixed-state (left panels), respectively the pure-state (right panels), reconstruction method. The top panels confirm the intuition that the gate error gradually increases with the system size for all dissipation strengths. This gradual shift is more pronounced for mixed-state reconstruction (top left panel of Fig. 3) than for the pure-state reconstruction (top right panel of Fig. 3).
@@
In the second scenario, only a single qubit is subject to dissipation, with rate $\gamma_{1}=\frac{1}{T_{1}}$ (bottom panels in Fig. 3). The dissipative qubit could be, e.g., a working qubit that is connected to memory qubits which do not exhibit the same level of noise [16]. Figure 3 (bottom) shows once again a gradual shift of the gate error with system size for the mixed-state reconstruction. This shift is smaller if only one qubits dissipates compared to the case where all qubits dissipate, cf. the top and bottom left panels of Fig. 3. The gate error of the pure-state reconstruction does not change appreciably if only one qubit dissipates, in contrast to the case where all qubits dissipate, cf. the top and bottom right panels of Fig. 3. Specifically, the gate error increases with $N$ visibly when all qubits are subject to dissipation but remains almost constant for single qubit decay.
@@
Since the differences of the gate error are relatively small on the scale used in Fig. 3, we analyze this behavior in greater detail for selected fixed $T_{1}$ decay times, see Fig. 4 for mixed-state and Fig. 5 for pure-state reconstruction.
@@
Figure 3: Gate error $1-F_{G}(U_{0},U_{RC})$ averaged over $n=\{100,75,50,10\}$ random unitary operations of the mixed-state reconstruction method versus $T_{1}$ decay time for $N=\{2,3,4,5\}$ qubits (for example for three qubits, i.e, $N=3$, the average gate error of $n=75$ different random unitary operations is calculated). The dissipation on the additional qubits with respect to a single qubit is (left) identical to the coupling of the first qubit, respectively (right) set to zero, i.e., only one qubit exhibits dissipation.
@@
Figure 4: Behavior of gate error $1-F_{G}(U_{0},U_{RC})$ of the mixed-state reconstruction method for the three $T_{1}$ times indicated by grey vertical dashed lines in Fig. 3.
@@
Figure 4 confirms our observation from Fig. 3 that for the mixed-state reconstruction, the gate error increases as a function of system size $N$ for both choices of dissipation and all $T_{1}$ decay times considered. The stronger scaling of the gate error with the system size in the mixed-state reconstruction, which persists even if only one qubit is affected by the decay, can be explained by the increased likelihood of a degeneracy to occur in the image of the mixed input state $\rho_{B}$: When the dimension of the Hilbert space increases, the eigenvalues of $\rho_{B}$ move closer together since $\text{Tr}[\rho_{B}]=1$. Moreover, even if no degeneracies occur, the ordering of the eigenvalues of the output states is more susceptible to errors for larger system sizes.
@@
For the pure-state reconstruction, the left panel Fig. 5 shows that if all qubits dissipate, the gate error increases with system size $N$. Conversely, if only one qubit dissipates, cf. the right panel in Fig. 5, the gate error of the pure-state reconstruction is roughly independent of the system size, with small fluctuations explained by the randomness in drawing the unitarities to be reconstructed in our simulations. This highlights that the consistently stronger performance of the pure-state reconstruction compared to the mixed-state reconstruction, observed in Sec. IV.3, persists and becomes more pronounced for larger system sizes.

### v3

@@ Abstract
We consider the problem of reconstructing, from a set of input and output states, the unitary describing the evolution of a quantum system or quantum channel. For ideal, fully coherent evolution, we show that the unitary can be reconstructed from two mixed states or $d+1$ pure states, where $d$ is the size of Hilbert space. The reconstruction method can be extended to approximate the unitary part of a dynamical map, provided the decoherence is not too strong to render this question meaningless. We exemplify the method for the example of the cross-resonance gate as well as a random set of unitaries, comparing the reconstruction from pure, respectively mixed, states to an approach based on the Choi matrix. We find that the pure state reconstruction requires the least amount of resources when the dynamics is close to unitary, whereas the mixed state approach outperforms the pure state reconstruction in terms of channel uses for appreciable decoherence. These conclusions hold also in the presence of SPAM errors and irrespective of the Hilbert space size.
@@ I Introduction
Here, we propose an alternative approach. The setup we envision is as follows: Considering an unknown or badly characterized quantum device, which input states should be prepared, sent through the channel, and measured on output, in order to minimize the number of channel uses? For an ideal, i.e., unitary, channel, the task of identifying the unitary can be thought of as two subtasks — first, to identify the basis in which the unitary is diagonal and then to determine the eigenvalues via their phases. The minimum number of states is therefore at least two, one that is built up of one-dimensional projectors, weighted with mutually disjoint eigenvalues, allowing to assess the information on the basis, and a second one for the phases, that is rotated with respect to all the one-dimensional projectors [5]. The second state can be thought of as a basis state taken from a mutually unbiased basis. The two states can be used to estimate the average fidelity of a quantum gate [5, 6], where one compares to a known (and desired) unitary evolution. Here we consider the more challenging task of reconstructing an unknown operation and show that measuring these two states upon output allows to reconstruct the unitary part of an unknown channel, provided that the decoherence is not too strong. The most obvious application is in channel learning. It is also useful when implementing quantum gates in architectures with limited connectivity where a series of auxiliary operations connecting the targeted qubits may distort the desired evolution in a non-trivial way. Identifying which unitary has been implemented instead of just knowing the error of the desired unitary will allow for designing tailored error mitigation strategies [3].
Since mixed states are hard to prepare, we also discuss a variant employing $d$ one-dimensional projectors instead of encoding the basis via one mixed state. While this ensures that all input states are pure, it takes their number from two (irrespective of Hilbert space dimension) to $d+1$, with $d$ the Hilbert space dimension. To benchmark the two state-based reconstruction approaches that we propose, we compare to reconstruction using the Choi matrix, i.e., full quantum process tomography.
The paper is organized as follows: In Sec. II we present our state-based reconstruction approaches for a unitary quantum channel using the minimal state set involving mixed, respectively pure, states. The generalization to noisy quantum channels is discussed in Sec. III, where we also outline the reconstruction based on the Choi matrix, as a reference point for our later analysis. In Sec. IV we apply the reconstruction algorithms via numerical simulations to a physical model involving typical relaxation processes. We start with two-qubit examples and then investigate the extension to multiple qubits. In Sec. V we compare the practical resource cost of the reconstruction algorithms in the context of state-of-the-art compressed sensing approaches, finding that it depends on whether the channel is close to unitary or experiences appreciable noise. Section VI summarizes our findings.
@@ II Reconstruction Approaches for Unitary Dynamical Maps
The action of a dynamical map on only two density matrices is sufficient to distinguish any two unitaries or to assess whether a given dynamical map is unitary [5]. Alternatively, it is possible, to employ a set of $d+1$ pure states, where $d$ is the dimension of the underlying Hilbert space [5]. This can be exploited to assess how well a quantum device realizes a desired prespecified operation with a minimal number of input states [5]. We now show that the results of Ref. [5] can be extended to reconstruct an unknown unitary operation $U$. Specifically, we demonstrate that the action of a dynamical map on at least two density matrices, respectively $d+1$ pure states, is sufficient to reconstruct the operation $U$ describing the action of a unitary dynamical map $\mathcal{D}_{U}(\rho)=U\rho U^{\dagger}$. We explain how $U$ can be obtained from the action of the dynamical map on the two minimal state sets mentioned above, making no further assumptions on the dynamical map except for its unitarity. This sets the stage for Sec. III where we extend the reconstruction to non-unitary dynamical maps 111while we will make no specific assumptions on the dynamical map, the question of extracting its unitary part is meaningful only for dynamical maps sufficiently close to unitary..
@@ IV Application
Next we test and compare the reconstruction methods outlined in Sec. III. We begin by introducing the physical model in Sec. IV.1 and introduce the reconstruction error $1-F_{G}$ in Sec. IV.2 as the main figure of merit to quantify the performance of the reconstruction methods. After analyzing the case of two qubits in Sec. IV.3, we investigate how the performance of the methods scales with the number of qubits in Sec. IV.4.
@@ IV.2 Figure of merit
The variance in Eq. (19) is taken over unitary maps with corresponding Choi matrix $\chi^{k}_{CL}$ corresponding to a Clifford gate $U^{k}_{CL}$ randomly drawn from the Clifford group. We use 100 random Clifford gates to evaluate the variance of their average fidelites with respect to the Choi matrix.
@@ IV.3 Results: Two qubits
Figure 1: Reconstruction error $1-F_{G}(U_{0},U_{RC})$ versus $T_{1}$ (left y-axis) of the reconstruction methods for random two qubit gates (left) and a CNOT gate (right). The shaded areas (left) represent the $1\sigma$ deviation with respect to the set of 100 random unitary gates drawn according to the Haar measure. The plot also shows the unitarity error $1-u(\chi)$ (right y-axis). The dissipation on the qubits is given by $L_{1}=\frac{1}{T_{1}}\sigma^{-}\otimes\mathds{1}$ and $L_{2}=\frac{1}{T_{1}}\mathds{1}\otimes\sigma^{-}$. The dissipation strength $T_{1}$ is given in units of the gate time $T_{g}$.
Figure 1 shows the reconstruction error $1-F_{G}$ for the mixed-state, pure-state, and Choi reconstruction method versus the $T_{1}$ decay time for random two-qubit operations (top panel) as well as the CNOT gate (bottom panel). It also shows the unitarity error $1-u(\chi)$ of the Choi matrix $\chi$ corresponding to the full dynamical map $\mathcal{D}(\cdot)$ according to Eq. (11). Note that the dissipation strength $\gamma_{1}$ of the dynamical map increases from right to left on the x-axis due to $\gamma_{1}=\frac{1}{T_{1}}$. Figure 1 clearly demonstrates that the pure-state and mixed-state reconstruction methods can be employed for dynamics that are sufficiently close to unitary, cf. the unitarity error (red line). Even when the decay time $T_{1}$ decreases, i.e. the dissipation strength increases, and the unitarity error of the dynamical map becomes larger, the two methods are able to approximately reconstruct the unitary part of the dynamical map for a wide range of dissipation strengths.
The lines for the reconstruction error of the mixed-state and pure-state reconstruction do not cover the entire range of $T_{1}$ times. This is due to the breakdown of the methods for large deviations from unitarity: When the dissipation strength increases, the assumptions used in Sec. II are no longer fulfilled, and degeneracies in the spectrum of the images $\mathcal{D(\rho)}$ prohibit the use of the pure-state and mixed-state reconstruction methods. In contrast, the reconstruction via the Choi matrix is not influenced by degeneracies in the image. Hence the corresponding reconstruction errors stretch across the entire range of $T_{1}$ times. Figure 1 furthermore shows that, for the mixed-state reconstruction, degeneracies in the image of the dynamical map occur for smaller dissipation strengths compared to pure-state reconstruction (compare the $T_{1}$-coverage of the orange and blue lines in Fig. 1). This is because the pure-state images in Eq. (7) are more protected against degeneracies than the mixed-state image (see Eq. (3)). It reflects the fact that for pure initial states it is less likely that the dissipative part of the dynamical map annuls the spectral gap between eigenvalues of different eigenvectors than it is for mixed initial states.
@@
We observe in Fig. 1 that the mixed-state reconstruction performs worse in terms of the reconstruction error $1-F_{G}$ than the pure-state reconstruction, while the Choi reconstruction performs best. This behavior is intuitive if one considers the total amount of information available: The Choi matrix contains, via Eq. (9), the information of the action of the dynamical map on arbitrary input states. The total amount of information accessible to the algorithm for the reconstruction is therefore higher compared to mixed-state and pure-state reconstruction. Similarly, since for mixed-state reconstruction a smaller number of states is propagated, less information is gained compared to the pure-state reconstruction. This explains the difference in reconstruction errors between mixed-state and pure-state reconstruction.
@@
For the mixed-state reconstruction of random gates, there is no appreciable difference in the general behavior as a function of $T_{1}$ and $T_{2}$, cf. the top panels of Fig. 1 and Fig. 2. In particular, larger reconstruction errors compared to the pure-state reconstruction method are observed for both dissipation processes. For the CNOT gates, two observations can be made: The reconstruction errors for both mixed-and pure-state reconstruction are larger than for the random gates. The dissipation strength, at which the image becomes degenerate, is larger for pure dephasing compared to $T_{1}$ decay (cf. the ranges of the blue and orange lines in the lower panels of Fig. 1 and Fig. 2). This is in contrast to the random gates, where there is no significant difference between $T_{1}$ and $T_{2}$ for both mixed-state and pure-state approach.
@@ IV.4 Results: Multiple qubits
How does the system size affect the performance of the reconstruction methods? To answer this question, we consider random unitaries on a system of $N$ qubits for the case of $T_{1}$ decay. We have seen in Sec. IV.3 that the reconstruction error of the reconstruction methods is connected to the unitarity of the dynamical map. The specific form of the dissipation determines unitarity, and we consider two choices in the following. In the first choice, all qubits are subject to decay, with identical dissipation rates $\gamma_{1}^{(i)}=\frac{1}{T_{1}}$ for $i=1,\dots,N$. Such a scenario is common in qubit arrays where the error is caused by spontaneous decay [17]. In this case, one would expect the reconstruction error to increase with system size, since the total dissipation strength is $\gamma_{1}^{\text{tot}}=\sum_{i}^{N}\gamma_{1}^{(i)}$ which scales with $N$. Figure 3 (top) shows the reconstruction error versus $T_{1}$ times for random unitaries for $N=2,3,4,5$ qubits, reconstructed using the mixed-state (left panels), respectively the pure-state (right panels), reconstruction method. The top panels confirm the intuition that the reconstruction error gradually increases with the system size for all dissipation strengths. This gradual shift is more pronounced for mixed-state reconstruction (top left panel of Fig. 3) than for the pure-state reconstruction (top right panel of Fig. 3).
@@
In the second scenario, only a single qubit is subject to dissipation, with rate $\gamma_{1}=\frac{1}{T_{1}}$ (bottom panels in Fig. 3). The dissipative qubit could be, e.g., a working qubit that is connected to memory qubits which do not exhibit the same level of noise [17]. Figure 3 (bottom) shows once again a gradual shift of the reconstruction error with system size for the mixed-state reconstruction. This shift is smaller if only one qubits dissipates compared to the case where all qubits dissipate, cf. the top and bottom left panels of Fig. 3. The reconstruction error of the pure-state reconstruction does not change appreciably if only one qubit dissipates, in contrast to the case where all qubits dissipate, cf. the top and bottom right panels of Fig. 3. Specifically, the reconstruction error increases with $N$ visibly when all qubits are subject to dissipation but remains almost constant for single qubit decay.
@@
Since the differences of the reconstruction error are relatively small on the scale used in Fig. 3, we analyze this behavior in greater detail for selected fixed $T_{1}$ decay times, see Fig. 4 for mixed-state and Fig. 5 for pure-state reconstruction.
@@
Figure 3: Reconstruction error $1-F_{G}(U_{0},U_{RC})$ averaged over $n=\{100,75,50,10\}$ random unitary operations of the mixed-state reconstruction method versus $T_{1}$ decay time for $N=\{2,3,4,5\}$ qubits (for example for three qubits, i.e, $N=3$, the average reconstruction error of $n=75$ different random unitary operations is calculated). The dissipation on the additional qubits with respect to a single qubit is (left) identical to the coupling of the first qubit, respectively (right) set to zero, i.e., only one qubit exhibits dissipation.
@@
Figure 4: Behavior of reconstruction error $1-F_{G}(U_{0},U_{RC})$ of the mixed-state reconstruction method for the three $T_{1}$ times indicated by grey vertical dashed lines in Fig. 3.
@@
Figure 4 confirms our observation from Fig. 3 that for the mixed-state reconstruction, the reconstruction error increases as a function of system size $N$ for both choices of dissipation and all $T_{1}$ decay times considered. The stronger scaling of the reconstruction error with the system size in the mixed-state reconstruction, which persists even if only one qubit is affected by the decay, can be explained by the increased likelihood of a degeneracy to occur in the image of the mixed input state $\rho_{B}$: When the dimension of the Hilbert space increases, the eigenvalues of $\rho_{B}$ move closer together since $\text{Tr}[\rho_{B}]=1$. Moreover, even if no degeneracies occur, the ordering of the eigenvalues of the output states is more susceptible to errors for larger system sizes.
@@
For the pure-state reconstruction, the left panel Fig. 5 shows that if all qubits dissipate, the reconstruction error increases with system size $N$. Conversely, if only one qubit dissipates, cf. the right panel in Fig. 5, the reconstruction error of the pure-state reconstruction is roughly independent of the system size, with small fluctuations explained by the randomness in drawing the unitarities to be reconstructed in our simulations. This highlights that the consistently stronger performance of the pure-state reconstruction compared to the mixed-state reconstruction, observed in Sec. IV.3, persists and becomes more pronounced for larger system sizes.
@@
IV.5 Results: Influence of SPAM errors
So far, we have assumed that all input states are prepared perfectly and that no read-out errors occur. We now investigate the influence of state preparation and measurement errors (SPAM) on the performance of the reconstruction methods. To this end, we adopt the model describing SPAM errors as unwanted bit-flip operations on both input- and output states [18, 19]. This yields the map
$\Phi[\rho]=(1-p)\rho+p\sigma^{x}\rho\sigma^{x}\,$, (20)
where $p$ denotes the probability that a SPAM error occurs. We apply $\Phi$ to the input states in Eqs. (1), (2) and (6) to account for state preparation errors. Similarly, to model measurement errors, we apply the same map to all output states of the dynamical map (e.g. $\mathcal{D}(\rho_{B})$ for the mixed intial state $\rho_{B}$).
Figure 6: Accounting for SPAM errors: Reconstruction error as a function of the SPAM probability $p$, cf. Eq. (20). for random two-qubit gates and amplitude damping with $T_{1}=7T_{g}$.
Figure 6 shows how the reconstruction error changes for increasing SPAM error probability $p$, assuming a moderate dissipation strength of $T_{1}=7T_{g}$. For $p=0$, the reconstruction error in Fig. 6 is the same as in Fig. 1 at the respective $T_{1}$ value. The moderate increase of the reconstruction error by at most one order of magnitude at the comparatively high SPAM error probability of 5% indicates an overall robustness of all three methods to SPAM errors. Perhaps unsurprisingly, pure-state and Choi reconstruction are somewhat more sensitive to SPAM errors than the mixed-state reconstruction. Importantly, for SPAM errors up to 1%, errors in the estimated output states do not seem to propagate to errors in the reconstructed unitary approximation, as the increase in the reconstruction error in this range is essentially negligible for all three methods in Fig. 6.